OpenStudy (anonymous):

Solve the following system using the substitution method. 2x-4y=0 y=x @eashy

OpenStudy (anonymous):

dO i put NO SOLUTION for this one?

OpenStudy (anonymous):

2x-4x=0 -2x=0 x=0

OpenStudy (anonymous):

Oooooh lol